# Breaking the Orthogonality Barrier in Quantum LDPC Codes

### v2

@@ 6.1 Structure and active orthogonality
Thus active orthogonality is equivalent to
@@
so the $r=3$ interaction can be kept free. The mother-matrix product $\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}$ is
@@ 6.2 Design goal and commutation table
With $L/2=6$ and $J=3$, we have $\Delta=\{0,1,2,4,5\}$ (so $3\notin\Delta$). The design goal is
$\Psi_{r}=0\ (r\in\Delta),\qquad\Psi_{3}\neq 0.$ (18)
This guarantees active orthogonality while breaking mother orthogonality, thereby creating room to prevent the latent part from becoming orthogonal to the active part.
As a schematic commutation table, suppose all pairs $(F_{i},G_{j})$ commute except $(i,j)=(0,3),(1,2)$. The table below shades commuting pairs in $\Gamma$ with 1:
@@
Then
$\Psi_{r}=0\ (r\neq 3),\qquad\Psi_{3}=F_{0}G_{3}+G_{3}F_{0}+F_{1}G_{2}+G_{2}F_{1}.$ (20)
A reader might think it is sufficient to keep only $\Psi_{3}$ nonzero while setting $\Psi_{4}=0$. We initially considered this and designed the code so that only $F_{0},G_{3}$ were non-commuting, yielding $\Psi_{3}\neq 0$ and $\Psi_{4}=0$. However, this produced many elementary trapping sets (ETS) formed by connecting $k$ length-8 cycles, which are the most harmful [27]. Representative ETSs are shown in Fig. 1. When we instead designed the code so that $\Psi_{3}\neq 0$ and $\Psi_{4}\neq 0$, making both $F_{0},G_{3}$ and $F_{1},G_{2}$ non-commuting, we substantially reduced connected ETSs for $k\leq 10$.
@@ 6.3 Affine permutation construction and choice of $P$
The affine permutations $f_{i},g_{i}$ constructed by the algorithm are listed in Table 1.
@@ 5 $373x+246$ $449x+672$
If $P$ is a prime power, there is no set of permutations $\sigma_{0},\sigma_{1},\tau_{0},\tau_{1}$ on $[P]$ that simultaneously satisfy: $\sigma_{0}$ commutes with $\tau_{0}$ but not with $\tau_{1}$, and $\sigma_{1}$ does not commute with $\tau_{0}$ but commutes with $\tau_{1}$. Using the standard semidirect-product representation of $\mathrm{AGL}_{1}(\mathbb{Z}/p^{k}\mathbb{Z})$, $(a,b):x\mapsto ax+b$, and the quadratic commutation condition, one can show this by analyzing the resulting zero-product conditions with $p$-adic valuation, which contradict non-commutativity. Therefore we choose $P=768=3\times 2^{8}$.
6.4 ETS library construction
In our initial experiments, ETSs of the form shown in Fig. 1 were the dominant cause of BP stalls. Since the constructed code has girth 8, we first enumerate 8-cycles and then build an ETS library by connecting 8-cycles to generate ETSs isomorphic to these patterns. The library is not a collection of only the ETSs encountered in decoding; instead, we enumerate all ETSs isomorphic to the dominant patterns and add them exhaustively.
Figure 1(a) is a $(6,2)$ ETS obtained by connecting two length-8 cycles, and Fig. 1(b) is a $(12,2)$ ETS obtained by connecting five length-8 cycles. Figure 2 shows a $(8,2)$ ETS formed by attaching a length-4 check–variable path between the two odd checks of a $(6,2)$ ETS. For the constructed code with $P=768,J=3,L=12$, the enumerator found $(6,2)$ ETSs: $X=48$, $Z=16$ (total 64); $(12,2)$ ETSs: $X=23$, $Z=0$ (total 23); and $(8,2)$ ETSs of the path4 type: $X=48$, $Z=0$ (total 48).
The parameters $f_{i},g_{i}$ constructed by the algorithm are listed in Table 1 (indices start at 0). Each $f_{i},g_{i}$ is an affine permutation $x\mapsto ax+b$ on $\mathbb{Z}_{P}$ with $\gcd(a,P)=1$. The resulting Tanner graph has girth 8 (no 4- or 6-cycles). We also verified that small $(a,2)$ ETSs obtained by connecting length-8 cycles do not appear up to a prescribed search limit $a\leq A$ (details omitted).
(a) $(6,2)$ ETS (two connected 8-cycles)
(b) $(12,2)$ ETS (five connected 8-cycles)
Figure 1: ETSs formed by connecting $k=2,5$ length-8 cycles. Filled check nodes have degree 1 (unsatisfied). Counts for $P=768,J=3,L=12$: $(6,2)$ ETS (X=48, Z=16) and $(12,2)$ ETS (X=23, Z=0).
Figure 2: Tanner graph of an ETS with (8,2). The internal subgraph matches the $(6,2)$ ETS in Fig. 1(a), and the length-4 path connects its former degree-1 checks (C0:599 and C0:542). Count for $P=768,J=3,L=12$: $(8,2)$ ETS (X=48, Z=0).
@@ 6.5 Minimum-distance evaluation
Commutativity makes the terms for $u=2,3,4,5$ vanish, so
@@
Thus $\Psi_{3}$ is a sum of four permutation matrices with disjoint supports, so both row and column weights are 4. Therefore any codeword $\boldsymbol{u}$ of $H_{Z}(\tilde{H}_{X})^{\mathsf{T}}$ decomposes as $\boldsymbol{u}=(\boldsymbol{u}_{3},\boldsymbol{u}_{4},\boldsymbol{u}_{5})$ where each $\boldsymbol{u}_{i}$ is a codeword of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$. From the structure of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$ and $\mathrm{Ker}(\Psi_{3})$, the support splits and weights are multiples of 4. The minimum distance of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$ is 4. We confirmed the existence of a weight-48 $\boldsymbol{x}$ constructed from a weight-4 codeword of $\Psi_{3}^{\mathsf{T}}$. Therefore $d_{X}^{(\mathrm{lat})},d_{Z}^{(\mathrm{lat})}\leq 48$, and consequently $d_{\min}\leq 48$. We do not prove a matching lower bound. Randomized searches did not find smaller logical errors, and large-scale decoding experiments did not exhibit logical failures, so we conjecture that this upper bound is tight.
6.6 FER performance
@@
Figure 3: FER curve of the constructed code girth-8, (3,12)-regular $9216,4612,\leq 48]]$, ($P=768,J=3,L=12$) with error bars (95% confidence). Here $n=PL$, the active matrix ranks are $\mathrm{rank}(H_{X})=2302$ and $\mathrm{rank}(H_{Z})=2302$, hence $k=4612$.

### v3

@@ 6.1 Structure and active orthogonality
Here we set $r:=(k-i)\bmod\,6$. For the active part we have $i,k\in\{0,1,2\}$, hence $r\in\Delta=\{0,1,2,4,5\}$. Thus active orthogonality is equivalent to
@@
so the $r=3$ interaction can be kept free. The parent-matrix product $\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}$ is
@@ 6.2 Design goal and commutation table
With $L/2=6$ and $J=3$, we have $\Delta=\{0,1,2,4,5\}$, so that $3\notin\Delta$. Our design guideline is to enlarge $d^{(\mathrm{lat})}$ as a proxy for increasing $d_{\min}$. In this setting, we target
$\Psi_{r}=0\quad(r\in\Delta),\qquad\Psi_{3}\neq 0.$ (16)
This condition guarantees active orthogonality while deliberately relaxing parent orthogonality, thereby creating room to prevent the latent part from becoming orthogonal to the active part.
@@
[…]
Under this commutation pattern, we obtain
$\Psi_{r}=0\quad(r\neq 3),\qquad\Psi_{3}=F_{0}G_{3}+G_{3}F_{0}+F_{1}G_{2}+G_{2}F_{1}.$ (18)
At first glance, one might expect that it is sufficient to keep $\Psi_{3}$ nonzero by making only $F_{0}$ and $G_{3}$ non-commuting. Indeed, we initially considered this minimal design and constructed codes satisfying $\Psi_{3}\neq 0$. However, this choice resulted in many $(a,b=2)$ elementary trapping sets formed by connecting multiple length-8 cycles, which are among the most harmful structures for BP decoding [28]. Moreover, two trapping sets sharing these same check-node pairs gave rise to low-weight logical operators. Representative examples are shown in Fig. 1. In contrast, by allowing both $(F_{0},G_{3})$ and $(F_{1},G_{2})$ to be non-commuting, we introduce additional degrees of freedom in $\Psi_{3}$, which substantially reduces the number of harmful trapping sets while still satisfying the active orthogonality constraint. In example codes constructed this way, the only small trapping sets of weight $a$ are those listed in Section 6.5, and none of them produced low-weight logical operators.
Remark.
Throughout Section 3 we considered the standard active choice consisting of the top $J$ block rows. More generally, let $S\subseteq\mathbb{Z}_{L/2}$ be an arbitrary active index set with $|S|=J$, and define
$\Delta_{S}:=\{(k-i)\bmod(L/2)\mid i,k\in S\}.$ (19)
The proof of Theorem 3.1 applies verbatim with $\Delta$ replaced by $\Delta_{S}$.
This generalization allows a substantial reduction of $|\Delta|$. In particular, for $J=3$ and $L=12$, choosing $S=\{0,2,4\}$ gives $|\Delta_{S}|=3$, whereas the standard choice yields $|\Delta|=5$. Such a reduction increases the number of indices $r\notin\Delta_{S}$ for which $\Psi_{r}$ can be kept nonzero, and may be exploited to further control trapping-set structures. Under this relaxed constraint, constructing codes may allow us to achieve the girth and minimum distance studied here at shorter codelengths.
@@ 6.3 Affine permutation construction and choice of $P$
We construct a girth-8 $(3,12)$-regular code; the parameters $f_{i},g_{i}$ are listed in Table 1 (indices start at 0). Each $f_{i},g_{i}$ is an affine permutation $x\mapsto ax+b$ on $\mathbb{Z}_{P}$ with $\gcd(a,P)=1$. The resulting Tanner graph has girth 8 (no 4- or 6-cycles).
@@ 5 $373x+246$ $449x+672$
If $P$ is a prime power, there is no set of affine permutations $\sigma_{0},\sigma_{1},\tau_{0},\tau_{1}$ on $[P]$ that simultaneously satisfy: $\sigma_{0}$ commutes with $\tau_{0}$ but not with $\tau_{1}$, and $\sigma_{1}$ does not commute with $\tau_{0}$ but commutes with $\tau_{1}$. Using the standard semidirect-product representation of $\mathrm{AGL}_{1}(\mathbb{Z}/p^{k}\mathbb{Z})$, $(a,b):x\mapsto ax+b$, and the quadratic commutation condition, one can show this by analyzing the resulting zero-product conditions with $p$-adic valuation, which contradict non-commutativity. Therefore we choose $P=768=3\times 2^{8}$.
6.4 Short Cycles
We begin by presenting a concrete example illustrating Theorem 3.4. For the $J=3,L=12$, the block positions
$W=F_{0}G_{0}^{-1}G_{5}F_{0}^{-1}F_{5}G_{5}^{-1}G_{0}F_{5}^{-1}.$
Counting such active block 8-cycles is also straightforward. For the active rows $0,1,2$, the unavoidable pattern uses adjacent left columns $i,i+1$ and adjacent right columns $j,j+1$ (indices modulo $L/2$), and it occurs when $i+j,i+j-1,i+j-2\in\Delta$. Thus the number of block 8-cycles in the active matrices is $(L/2)|\Delta\cap(\Delta+1)\cap(\Delta+2)|$. For $J=3,L/2=6$, $\Delta=\{0,1,2,4,5\}$ and $\Delta\cap(\Delta+1)\cap(\Delta+2)=\{0,1,2\}$, so there are 18 block 8-cycles, yielding $18P$ lifted 8-cycles.
Enumerating cycles in the active Tanner graphs of this instance gives 60,51260{,}512 8-cycles on the $X$ side, and 54,65654{,}656 8-cycles on the $Z$ side. By Theorem 3.4, 8-cycles are unavoidable in the active Tanner graphs; the block pattern above yields the lower bound 18P=13,82418P=13{,}824. The total exceeds this bound because commutativity holds for many additional pairs and the specific APM parameters create extra cycles beyond block lifts, so the total need not be a multiple of $P$.
6.5 Trapping set library construction
In our initial experiments, trapping sets of the form shown in Fig. 1 were the dominant cause of BP stalls. Since the constructed code has girth 8, we first enumerate 8-cycles and then build an trapping set library by connecting several 8-cycles to generate trapping setes isomorphic to these patterns. The library is not a collection of only the trapping sets encountered in decoding; instead, we enumerate all trapping sets isomorphic to the dominant patterns and add them exhaustively.
Figure 1 shows three elementary trapping-set patterns: $(6,2)$ from two length-8 cycles, $(12,2)$ from five length-8 cycles, and $(8,2)$ formed by attaching a length-4 check–variable path to the two odd checks of (6,2). For $P=768,J=3,L=12$, the counts were: $(6,2)$ X-side: 48, Z-side: 16 (total 64); $(12,2)$ X-side: 23, Z-side: 0 (total 23); and $(8,2)$ path4 X-side: 48, Z-side: 0 (total 48).
$(6,2)$ ETS (two connected 8-cycles)
$(12,2)$ ETS (five connected 8-cycles)
$(8,2)$ ETS (path4 attachment)
Figure 1: Representative ETSs used in the library.
@@ 6.6 Minimum-distance evaluation
Recall the latent-based distances $d_{X}^{(\mathrm{lat})}$ and $d_{Z}^{(\mathrm{lat})}$ defined in Section 2. We now construct explicit low-weight vectors in the latent row spaces that satisfy the active parity checks but are not in the duals. Equivalently, any $\boldsymbol{x}\in\operatorname{Row}(\tilde{H}_{X})$ can be written as $\boldsymbol{x}=(\tilde{H}_{X})^{\mathsf{T}}\boldsymbol{u}$ for some $\boldsymbol{u}$, and likewise for $Z$. Concretely, we look for coefficient vectors $\boldsymbol{u}\in\mathrm{Ker}\!\bigl(H_{Z}(\tilde{H}_{X})^{\mathsf{T}}\bigr)$ and $\boldsymbol{v}\in\mathrm{Ker}\!\bigl(H_{X}(\tilde{H}_{Z})^{\mathsf{T}}\bigr)$, then lift them to $\boldsymbol{x}=\boldsymbol{u}^{\mathsf{T}}\tilde{H}_{X}$ and $\boldsymbol{z}=\boldsymbol{v}^{\mathsf{T}}\tilde{H}_{Z}$. Any such $\boldsymbol{x},\boldsymbol{z}$ yields $d_{X}^{(\mathrm{lat})}\leq\operatorname{wt}(\boldsymbol{x})$ and $d_{Z}^{(\mathrm{lat})}\leq\operatorname{wt}(\boldsymbol{z})$, so below we exploit the block-diagonal form to obtain explicit weight-48 examples.
@@
Thus $\boldsymbol{u}=(\boldsymbol{u}_{3},\boldsymbol{u}_{4},\boldsymbol{u}_{5})$ and $\boldsymbol{v}=(\boldsymbol{v}_{3},\boldsymbol{v}_{4},\boldsymbol{v}_{5})$ with $\boldsymbol{u}_{r}\in\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$ and $\boldsymbol{v}_{r}\in\mathrm{Ker}(\Psi_{3})$ for $r=3,4,5$. Commutativity makes the terms for $u=2,3,4,5$ vanish, so
@@
In this instance, $\mathrm{rank}(\Psi_{3})=576$, so $\dim\mathrm{Ker}(\Psi_{3})=\dim\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})=192$. The four permutation terms in $\Psi_{3}$ have disjoint supports, so both row and column weights are $4$ and the kernels are generated by $192$ disjoint weight-$4$ codewords supported on the $4$-point blocks
$t]:=\{t,\ t+192,\ t+384,\ t+576\}\subset\mathbb{Z}_{768},\qquad t\in\mathbb{Z}_{192}.$ (24)
The explicit weight-$4$ vectors constructed above lift to $\boldsymbol{x}=\boldsymbol{u}^{\mathsf{T}}\tilde{H}_{X}$ and $\boldsymbol{z}=\boldsymbol{v}^{\mathsf{T}}\tilde{H}_{Z}$ of weight $48$, and they are not in $C_{X}^{\perp}$ and $C_{Z}^{\perp}$ for this instance. This gives the upper bounds $d_{X}^{(\mathrm{lat})}\leq 48$ and $d_{Z}^{(\mathrm{lat})}\leq 48$.
Because each APM block preserves the $4$-point blocks, any latent vector $\boldsymbol{x}=\boldsymbol{u}^{\mathsf{T}}\tilde{H}_{X}$ is block-constant within each column block. Compressing each column block to length $192$ gives $\bar{\boldsymbol{x}}$ with $\operatorname{wt}(\boldsymbol{x})=4\,\operatorname{wt}(\bar{\boldsymbol{x}})$. In the compressed space, each generator has weight $12$ and overlaps across row blocks are uniformly bounded, so any nonzero $\bar{\boldsymbol{x}}$ has $\operatorname{wt}(\bar{\boldsymbol{x}})\geq 12$. Hence $\operatorname{wt}(\boldsymbol{x})\geq 48$, and the same argument applies to $Z$, so $d_{X}^{(\mathrm{lat})}\geq 48$ and $d_{Z}^{(\mathrm{lat})}\geq 48$ are provable here. Together with the explicit weight-$48$ examples we conclude $d_{X}^{(\mathrm{lat})}=d_{Z}^{(\mathrm{lat})}=48$. This shows $d_{\min}\leq 48$.
To conclude $d_{\min}=48$, we must also bound logical operators outside the latent range; in other words, we need a lower bound on the minimum weight of $\boldsymbol{x}\in C_{Z}\setminus C_{X}^{\perp},\boldsymbol{x}\notin\operatorname{Row}(\tilde{H}_{X}).$ At present we do not have certified lower bounds on these non-latent minima, so the rigorous statement here is only $d_{\min}\leq 48$. In summary, we prove $d_{X}^{(\mathrm{lat})}=d_{Z}^{(\mathrm{lat})}=48$ for this instance, which yields the bound $d_{\min}\leq 48$, while certified lower bounds for non-latent logical operators remain open.
6.7 Frame error rate
@@
Figure 2: FER curve of the constructed code girth-8, (3,12)-regular $9216,4612,\leq 48]]$, ($P=768,J=3,L=12$) with error bars (95% confidence). For reference, we plot the depolarizing-channel hashing bound and the density-evolution benchmark threshold for a non-orthogonal random $(3,12)$-regular $H_{X},H_{Z}$ pair ($p=0.05702$). Here the number of physical qubits is $n=PL$, the active matrix ranks are $\mathrm{rank}(H_{X})=2302$ and $\mathrm{rank}(H_{Z})=2302$, hence the number of logical qubits is $k=4612$.
6.8 Density evolution benchmark
To benchmark BP performance, we analyze a non-orthogonal random $(3,12)$-regular pair of classical LDPC codes $H_{X},H_{Z}$ drawn independently from the standard configuration-model ensemble. Although this pair is no longer a CSS code and has no direct physical meaning, it provides a useful classical benchmark for estimating noise from syndromes. Under the cycle-free (infinite-length) approximation, we estimate the density-evolution fixed point using a population-dynamics (Monte Carlo) simulation: following the numerical viewpoint emphasized by MacKay [19], we track a large population of messages when analytic density tracking is intractable, iteratively updating it by randomly sampling tree-like neighborhoods with the regular degrees and applying the BP update rules for the depolarizing channel; see [26, 27] for standard density-evolution background. This yields a depolarizing-channel BP threshold of $p\simeq 0.05702$, which we plot in Fig. 2 as a benchmark reference. The proposed code approaches this DE threshold, suggesting it retains sufficient LDPC randomness and that the decoder effectively exploits it.
Data availability
The data that support the findings of this study are available from the author upon reasonable request.
Code availability
The construction algorithm is available at https://github.com/kasaikenta/construct_apm_css_code. The decoding algorithm is available at https://github.com/kasaikenta/joint_BP_plus_PP.